OpenStudy (anonymous):

evaluate r+s/6 for r = 15 and s = 15 simplify your answer r+s/6 =

myininaya (myininaya):

15+15/6=15(6/6)+15/6=90/6+15/6=105/6

OpenStudy (saifoo.khan):

35/2

myininaya (myininaya):

yes after simplifying 105/6 you get 35/2

OpenStudy (saifoo.khan):

(:

myininaya (myininaya):

you want to know a trick?

OpenStudy (anonymous):

yes

myininaya (myininaya):

if the sum of the digits of a number is divisble by 3, then that number is divisible by 3. like 105 1+0+5=6 6 is divisible by 3 so 105 is divisible by 3

OpenStudy (anonymous):

thanks both of you